Project

General

Profile

NDOS CommissionMtng 20120827

Attending Mat, Peter, Gavin, Andrew, Jan, Susan, Leon

UPDATES:
1.  Detector Update

14069 - DDS was restarted on master. Seen as errors in Run control and AppMan. Data taking continued but everything is n subrun 14.

Zukai performed extend DSO studies on Wednesday: Extend DSO 4 hours for pixels. Scan in high detail. Extend to more channel? Cooling states? (Requires new DCMProgUtils/DCMReadout backward compatible with firmware which currently breaks PedestalData Runner)

DSO speed gains. -Factor of 2 for standard

DCS work

2.  DataCheck Update

Jan- Aug 17 weekend ran though 14055/16
200k event in 14049
3-2 port 51 shut off in subrun 0 in 4/5 runs.

4.  Installation Plans for the week

-Parylene Coated APD through out the week…
-10-20 per day
-Actual detector

5.  DAQ Update and Plans for the week
NOVA_DAQ_DCS update will happen now.
Firmware still in the works.

6.  DCS Update and Plans for the week
Gennadiy and Steve asked how to update NOVADCS
Kurt want to setup new database table for dry air association.

7.  AOB
None